Prev Next

SW&IT Meeting Minutes 2018-09-24

Attendees: Andrew, David, Igor, Kellee, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

News from management

  • SOUL 1st light
  • s/w recruitment

Previous Action Items

  • Vsphere (DE)
  • power plan update draft (DE)
  • Luci H/O doc (IS)
  • Luci telemetry
    • web-io cold heads state for vibration (IS)
    • luci telemetry in git (DONE)
  • vm backup

Transition plan

Handover steps

  1. primary + secondary do self learning, gather information, read wiki
  2. primary organize teamviewer/zoom walk through with Kellee and secondary
  3. primary draft techtalk w. input from secondary
  4. techtalk review by Kellee
  5. primary does a team presentation

Please update this table to indicate progress.


Application(s) Primary Secondary WeekSorted ascending
OAC
AC (2)
Petr (2)
1
ALH Igor(1/2)
Matthieu (2)
1
GCS
Petr (3)
Matthieu (2)
1
AzCam
AC (2)
Petr (2)
2
DIMM
AC (2)
Petr (2)
2
Telemetry
Petr (3)
Matthieu (2)
2
Status Server Stephen (1)
Matthieu (1)
2
LBC
Yang (1/2)
Igor(1/2) 2
MODS
Yang (1)
Igor 2
Weatherstation
AC (2)
Petr 3
INDI Igor Matthieu 3
OVMS
Petr (3)
Matthieu (2)
3
Telemetry cleanup Stephen Petr 3
Telem-VIZ Yang Stephen 3
FACSUM,WEATHERSTATION
AC (2)
Matthieu 4
particle monitors Yang Petr 4

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

This week in github

Team reports

Team reports

PK

  • various patches (IT7380, IT7407, IT 7408, IT3949)
  • GCS changes, PCS changes
  • tpk documentation reading (primary for copoting and LBC NS guidigin issues)
  • git discover --tags

other:
  • sucessfully used "git bisect" to find cause of an odd bug (cause by adding setlocale call to one of the components) - if you have a smoking gun, easy to reproduce error, git bisect is a great tool
  • skyflats, few bug fixes on AzCam driver (Kuiper 61" over weekend)

next:
  • GCS split, readibility changes
  • LBC NS guiding

YZ

  • Getting started with MODS (but not much yet).
  • Schedule another meeting with Kellee on LBC/MODS discussion (tentative on Tuesday).
  • Focused on producing a revised collimation tool, as well as revamping previous telemetry data search tool.
  • Continue working on documenting pointing and collimation procedures.
  • Continue working on Python package and modules.
  • Need to start gather more info on Al's AO project.

MB

  • management OC, WPC
  • Labor Reporting, fixing odd cases
  • TMT ESW review, much more material than anticipated
  • attend SH LBTI/single mode fiber meeting
  • SHark NIR follow up ...
  • jenkins/docker work, recover issue
  • Twiki migration
  • obs workstation upgrade shared with OC (but skipped)
  • ims vm+ks+ci-rpm / with SH
  • LBR stats, preview

next:
  • EngMgr interviews Mon/Tue, 2/3 on site candidates
  • TMT ESW review Thu. full day
  • vm backup with DE
  • IT to github issues (overdue), IT reassignments
  • cisrv01 rebuilt, minimal plugins, git capture
  • IA2 ??
  • ALTA

AC

Last week:
  • Prepare OAC, AzCam, DIMM, FACSUM/Weather Station tech talks
  • Investigate MPIA repo inconsistencies
  • ARGOS deployment

This week:
  • All of the above
  • IT 6299 (FACSUM AO State)
  • IT 7388 GitHub review
  • IT 7421 (OAC telemetry)

KS

Last week:
  • Vacation Monday and Tuesday.
  • Looked into xenomai watchdog complaints on OVMS RACK from acquisition task - the main thread. UEI folks said we're just missing a sleep. I modified the code and it's ready for a test.
  • FACSUM/WEATHERSTATION discussion with AC.
  • John Hill suggested the latest OVMS software may have the wrong code for units or a wrong config file. Searched around in my old email and notes to see when things changed. I'm confident the units are correct in the source code, but I found that we have not been using the "no M1" config file for awhile.

This week:
  • Status Server meeting with SH and MB.
  • LBC/MODS discussion with YZ.
  • Review Petr's presentations - GCS, OVMS, telemetry

DE

Last week:
  • Finally got Peter's laptop
  • Discovered backup problems with ciserv
    • Probable solution found
  • Modified backup options in bulk backup
    • Over the weekend, I discovered a problem with this option and XP; will implement a fix for that
    • (Today) Installed VMWare tools on all Linux VMs to support; will likely need reboots on:
      • eap-controller

      • isola-luci

      • ksummers-dev

      • mbec-dev

      • osa-tcs

      • pkubanek-dev

      • pkubanek-dev1
      • xzhang-dev
      • yzhang-dev
This week:
  • Peter laptop setup
  • Found suboptimal configuration of mountain VM host after Friday's reboot--changed to automatically start VMs with host start
  • Test-restore recent backups with new backup options

IS

Last Week:
  • LUCI
    • Mostly LUCI RTD/AFC work
    • Release version done on Friday. Dave T. is testing
  • OT
    • LBC xsl bug fix
    • Creating a script to test otscripting for Zabbix (not done)
This Week:
  • PIT work
    • Fixing bug about time units and release.
  • LUCI RTD work
    • Fixing bug on default LMS loaded
    • Reading LMS info from FITS headers instead of LMS file downloaded from FTP
    • AFC
      • support Dave with AFC test work as needed
  • Documentation
    • Need to go though the LUCI handover and send email to Dave T. with questions.

SH

Last week:
  • LBTI fiber meeting
  • statserv gander
  • VM host problems on mountain
  • rm507 mirroring issue
  • cisrv01 changes

This week:
  • rebuild cisrv01
  • Foswiki
  • docker repository
  • fiber installation/networking help
  • power shutdown Tuesday.
Topic revision: r15 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback